Hazelnut Corp. manufactures lawn ornaments. It currently has two product lines, the basic and the luxury. Hazelnut has a total of $182,465 in overhead.

The company has identified the following information about its overhead activity cost pools and the two product lines:

Activity Cost Pools Cost Driver Cost Assigned
to Pool
Quantity/Amount Consumed by
Basic
Quantity/Amount Consumed by
Luxury
Materials handling Number of moves $ 4,200 22 moves 48 moves
Quality control Number of inspections $ 34,265 240 inspections 145 inspections
Machine maintenance Number of machine hours $ 144,000 4,800 machine hours 4,800 machine hours

Required:
1.
Suppose Hazelnut used a traditional costing system with machine hours as the cost driver. Determine the amount of overhead assigned to each product line. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount.)



2. Calculate the activity rates for each cost pool in Hazelnut’s ABC system.



3. Calculate the amount of overhead that Hazelnut will assign to the basic line if it uses an ABC system.



4. Determine the amount of overhead Hazelnut will assign to the luxury line if it uses an ABC system.



5. Compare the results for a traditional system with that of an ABC system. Which do you think is more accurate?

Traditional System
ABC System
Answer #1

1) Overhead assigned under traditional costing

Basic line Luxury line
Overhead assigned 182465/9600*4800 = 91233 91232

2) Activity rate

Material handling = 4200/70 = 60 per move

Quality control = 34265/385 = 89 per inspection

Machine maintenance = 144000/96000 = 15 per machine hour

3) Overhead assigned to Basic line

Basic line
Material handling 60*22 = 1320
Quality control 89*240 = 21360
Machine maintenance 72000
Total overhead assigned 94680

4) Overhead assigned to Luxury line

Luxury line
Material handling 60*48 = 2880
Quality control 89*145 = 12905
Machine maintenance 72000
Total overhead assigned 87785

5) ABC Cost